# PACE: Procedural Abstractions for Communicating Efficiently

### v2

@@ Abstract
Keywords: efficient communication; reinforcement learning; abstractions learning;
@@ Introduction
Procedural tasks such as cooking and programming require executing a sequence of actions to achieve a desired goal. A natural approach to reduce their complexity and improve generalisation to new tasks is to introduce abstractions for common sub-sequences [Solway et al., 2014]. For example, in cooking, techniques such as sautéing or kneading serve as foundational building blocks that simplify complex recipes. Similarly, abstractions emerge in repeated communication between human dyads when collaborating on shared tasks [Krauss and Weinheimer, 1964, Hawkins et al., 2020, McCarthy et al., 2021]. Over time, as new shared abstractions are introduced, the shared language becomes more concise. This improves cooperation and allows humans to reach their goals more effectively.
How communication shapes abstractions is of much interest within the AI and Cognitive Science communities [Lee, 1996, Ho et al., 2019]. A compelling hypothesis is Efficient Communication which argues that languages are under pressure to be informative whilst minimising cognitive load (e.g. vocabulary size) [Kemp and Regier, 2012, Gibson et al., 2017, Zaslavsky et al., 2019, Gibson et al., 2019]. This provides insights into the abstractions humans converge to in semantic domains such as colour, kinships and others [Xu et al., 2020, Regier et al., 2015, Kemp and Regier, 2012, Yin et al., 2024]. Denić and Szymanik [2024] demonstrated how in other domains, in which semantics are compositional and informativity can be maximal even with very few terms, a different trade-off shapes languages: a pressure to lexicalise as few concepts as possible while simultaneously keeping utterances as brief as possible (i.e. minimize the average morphosyntactic complexity). Within collaborative procedural tasks, which pressures for efficient communication impact abstraction introduction and use and in what way? We take the first steps to answer this question within the context of artificial agents.
@@
Figure 1: In early rounds of the architect-builder game, the architect messages refer to horizontal or vertical blocks (a). After multiple interactions, the architect tries to introduce an abstraction (b), which after a learning period allows for shorter communication to solve the task (c).
@@
The architect-builder game introduced in McCarthy et al. [2021] and shown in Figure 1 provides a natural domain to study this problem. It is a repeated game which features two-agents called the architect and the builder who must complete a collaborative building task. In each round (or iteration) the architect observes the goal-scene and conveys some instructions in natural language to the builder. The builder interprets these instructions and attempts to reconstruct the scene by placing blocks within the grid. If the goal is not reached, the architect may send further instructions. In human experiments, more informative procedural abstractions emerge over successive rounds as communication strategies are refined, leading to more effective cooperation. Although computational models of this process have been proposed, they rely on simplifying assumptions: McCarthy et al. [2021] explicitly constrain the number of abstractions the architect can introduce. They also model the builder with a predefined discrete mapping from instructions to shapes they should build, providing them with the right response to tokens that refer to shapes that may have not yet been introduced by the architect. Jergéus et al. [2022] instead assume that the builder has an immediate perfect understanding of the new abstractions introduced by the architect. Both of these approaches limit the ability to capture the collaborative dynamics of language learning in different ways, which we address in this work.
We propose a novel multi-agent neuro-symbolic method called Procedural Abstractions for Communicating Efficiently (PACE). By integrating library learning [Ellis et al., 2021, Bowers et al., 2023], emergent communication (EC) [Foerster et al., 2016, Lazaridou and Baroni, 2020], and reinforcement learning (RL) [Sutton and Barto, 1998], PACE offers a unified framework for studying the formation and evolution of abstractions across multi-round interactions. This approach allows for constructing interpretable procedural abstractions via symbolic library learning, while neural EC methods enable the development of a flexible, learnable communication language. In previous work, EC has been used to explain how communicative pressures for efficient communication shape the language structure in other settings [Carlsson et al., 2021, 2024]. We apply EC techniques to study how pressures manifest within abstraction learning. Whilst pure EC methods like LazImpa [Rita et al., 2022] have been used to learn efficient languages they are not readily extendable to the multi-interaction settings which we are interested in. Our setting is challenging as reward is sparse and the task occurs over a long time-horizon when compared to standard Lewis signalling games [Lewis, 1969]. PACE mitigates these issues and allows for easy study of these problems.
@@
We evaluate PACE on an extension of the Architect-Builder game. In our extension, the architect is a neuro-symbolic agent that composes programs to achieve the goal-scene within an internal symbolic language. The programs are neurally encoded as messages, following conventions from EC, and are then interpreted by the builder. After several rounds, the architect’s internal symbolic language is extended with new abstractions via a probabilistic library learning method. These new abstractions provide alternative ways to express goal-scenes allowing for more concise programs but require the agents to learn how to communicate and understand these new abstractions. This contention between shorter but unestablished programs and more verbose established programs is handled via RL methods. We find that it exhibits similar tendencies to humans — a development toward a richer language that allows for more concise utterances. Interestingly, after a number of abstractions have been introduced, our model naturally converges to a stable language. Moreover, we show that in this setting languages that are closer to optimality in terms of the trade-off between average morphosyntactic complexity and language size are easier to learn, connecting our work with the Efficient Communication literature. Our approach addresses limitations of existing approaches and provides a valuable framework for future exploration in this area.
@@
We extend the dataset from McCarthy et al. [2021] for our experiments. We increase the size going from $3$ to $31$ unique shapes, with multiple sub-shapes reoccurring in different shapes. Our shapes are of different sizes and resemble either uppercase or lowercase letters from the English alphabet (see Figure 10 in the Appendix). As before, the dataset consists of scenes composed of two shapes placed side by side. Our dataset contains $961$ goal-scenes (compared to the original $9$), which enables us to use it for training neural agents. We split the dataset into 930 training scenes and 31 test scenes. These splits are constructed to ensure the distribution over shapes is the same. However, in the test set, the ordered pair of shapes constituting each goal-scene, does not appear in the training set.
@@
There are three phases to PACE, Selection, Communication and Abstraction (see Figure 2). The architect must first choose a program to communicate. For this purpose the architect is initialised with a table of programs written in ${\cal A}_{init}$ for each scene in the training set. Initially there is one program per scene. Secondly, the architect learns how to communicate (neurally) instructions to the builder via EC, allowing them to learn a language for collaboration. After some rounds, the architect enters the Abstraction phase, where it uses a library learning mechanism to identify common sub-sequences to abstract. The new abstraction is then used to add additional (shorter) programs for the relevant scenes. In the next Communication phase, the architect thus has multiple programs to choose from. We use reinforcement learning techniques to learn to choose between alternative programs. We next describe these three phases in more detail.
@@
As in other settings like recursive numeral systems [Denić and Szymanik, 2024], programs in the Architect-Builder game can be perfectly informative even with a minimal vocabulary – precisely of size two, with one term referring to the horizontal primitive block and the other referring to the vertical one. With just the two initial action-words the architect can describe infinitely complex goal-scenes, given enough time! This is because the semantics of the environment are compositional. Thus, as Denić and Szymanik [2024] show for recursive numeral systems, we expect that the pressures in competition to shape the agents’ languages in the Architect-Builder game are the average morphosyntactic complexity of the programs describing the goal-scenes. In our case this corresponds to a trade-off between the average length of the programs communicated by the architect, and the number of lexicalised terms, i.e. the number of primitives plus abstractions in $\mathcal{A}$, the architect’s lexicon.
@@
Through empirical experimentation we want to investigate the following questions: 1) Does PACE display conversational tendencies which are similar to humans, allowing for more concise communication after multiple interactions? 2) Does communicative pressures impact what abstractions get adopted? If so, which abstractions are adopted and which are discarded? The behaviour of PACE is compared to two naïve baselines, to assess the impact of each of PACE components better. The first, No abstractions, is PACE without the abstraction phase. The second, Greedy, always picks the shortest program, eliminating the bandit. We use reward (also referred to as reconstruction accuracy) and average morphosyntactic complexity to compare these approaches. Unless otherwise stated results obtained are averaged over $16$ runs. We provide further implementation details in the Appendix.
@@
Figure 7: Two goal-scenes with representations of their initial and final programs chosen by ACE.
@@
McCarthy et al. [2021] propose a Bayesian model of procedural abstractions that combines library learning [Ellis et al., 2021] with social reasoning-based communication [Goodman and Frank, 2016]. While effective in capturing human-like trends, this model assumes a predefined mapping from instructions to actions the builder should take, essentially defining a priori the meaning of messages that the architect might not have introduced yet. They also explicitly constrain language size, adding a parameter that makes it impossible to add further abstractions after a certain number have been introduced. Jergéus et al. [2022] take a Deep Reinforcement Learning approach but oversimplify interactions by assuming the builder immediately understands abstractions, collapsing the setup into a single-agent framework and removing pressures for efficient communication. By using Emergent Communication (EC) within PACE, our approach enables both architect and builder to learn interactively, introducing abstractions organically and letting communicative pressures naturally limit language size.
@@ Conclusion
In collaborative task-orientated communication humans tend towards more concise utterances by introducing procedural abstractions. In this work we propose a novel neuro-symbolic algorithm called PACE which displays similar tendencies. PACE integrates ideas from library learning, emergent communication and reinforcement learning allowing for intepretable study of how abstractions are constructed and introduced in task-orientated communication. Our work serves as a bridge between procedural abstraction learning and efficient communication [Kemp and Regier, 2012, Gibson et al., 2017, Zaslavsky et al., 2019, Gibson et al., 2019, Denić and Szymanik, 2024, Carlsson et al., 2024]. We demonstrate that more optimal languages are easier to learn reinforcing ideas relating learnability with efficient communication.
In future work, we intend to extend our analysis to consider other domains where languages are perfectly informative and trade-off average morphosyntactic complexity and vocabulary size, for example, in recursive numeral systems. We also intend to explore how LLMs understand and reason about these conversational dynamics. Significant research effort is being invested into exploring the role of natural language as a mechanism for humans to provide instructions to intelligent agents [Brohan et al., 2023, Shi et al., 2024]. Providing these systems with the capability of handling novel procedural abstractions can facilitate improved cooperation between humans and agents. We believe that PACE represents a step towards equipping intelligent agents with flexible and extendable languages.
@@ Acknowledgments
This work was supported by funding from from the Wal- lenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foun- dation and the Swedish Research Council. The computa- tions in this work were enabled by resources provided by the Swedish National Infrastructure for Computing (SNIC). We also extend our thanks to Emil Carlsson for insightful discussions at various points throughout the project.
@@
Implementation Details
We implement PACE within Python version 3.9.18 and PyTorch version 2.2.1 [Paszke et al., 2019] and run experiments on Apple M2 Pro CPU (16 GB) on MacOS 14.4.1. The communication policies, $\pi_{comm}$ and $\pi_{bldr}$, are deep neural networks which have $1$ and $2$ hidden layers, respectively. Each hidden layer contains $200$ neurons and utilises ReLU activation functions. The outputted layer of $\pi_{comm}$ is comprised of $30$ neurons, which is its maximum vocabulary size. The output layer of $\pi_{bldr}$ is comprised of $81$ neurons, which is the size of the grid. Each output neuron in $\pi_{bldr}$ represents the builder’s belief that the grid index is occupied. We use the gumbel-softmax relaxation with the hard parameter so that the networks can be differentiated end-to-end. The model is optimised to minimise binary cross entropy error and we use a learning rate of $0.0009$ in conjunction with the ADAM optimiser [Kingma and Ba, 2014]. The bandit hyperparameters $\alpha$, $\gamma$, $\epsilon$ and $q_{init}$ are set to $0.5$, $0.99$, $0.1$, and $0.0$ respectively. Note, that before an abstraction is introduced we prune all but the $3$ best programs (determined empirically) for each goal-scene in $\pi_{arch}$ to not incur an exponential growth in the number of programs for each goal-scene. We also set the number of epochs $e$ to 40. All hyperparameters chosen are experimentally determined through grid search.

### v3

@@ Abstract
Keywords: efficient communication; reinforcement learning; abstractions learning.
footnotetext: Correspondence to: silvi@chalmers.se.
@@ Introduction
Procedural tasks such as cooking and programming require executing a sequence of actions to achieve a desired goal. A natural approach to reduce their complexity and improve generalisation to new tasks is to introduce abstractions for common sequences of actions [Solway et al., 2014]. For example, in cooking, techniques such as sautéing or kneading serve as foundational building blocks that simplify complex recipes. Similarly, abstractions emerge in repeated communication between human dyads when collaborating on shared tasks [Krauss and Weinheimer, 1964, Hawkins et al., 2020, McCarthy et al., 2021]. Over time, as new abstractions are introduced into the shared language, communication becomes more concise. This can improve cooperation and allow to reach the goal more easily.
@@
Figure 1: Two artificial agents playing the architect-builder game, starting from a small artificial language. Initially, the architect messages refer to horizontal or vertical blocks (a). After multiple interactions, the architect tries to introduce an abstraction (b), which after a learning period allows for shorter communication to solve the task (c).
How communication shapes abstractions is of much interest within the AI and Cognitive Science communities [Lee, 1996, Ho et al., 2019]. A compelling principle is Efficient Communication, which argues that languages are under pressure to be informative whilst minimising cognitive load [Kemp and Regier, 2012, Gibson et al., 2017, Zaslavsky et al., 2019, Gibson et al., 2019]. This provides insights into the abstractions humans converge to in semantic domains such as colour, kinships and others [Xu et al., 2020, Regier et al., 2015, Kemp and Regier, 2012, Yin et al., 2024]. Here we want to explore the principle of Efficient Communication with artificial agents in collaborative procedural tasks: how do the pressures for efficient communication manifest and impact abstraction introduction and use?
@@
The architect-builder game introduced in McCarthy et al. [2021] provides a simple framework to study the use of abstractions in collaborative tasks. It is a repeated game with two participants, the architect and the builder, who must complete a collaborative building task. In each round, the architect observes the goal-scene on their screen and conveys some instructions in natural language to the builder. The builder, who cannot see the original scene, interprets these instructions and attempts to reconstruct the shapes by placing blocks within the grid. The architect may send further instructions to help the builder. McCarthy et al. [2021] showed that in early rounds human participants use lengthy utterances, while later realizing it is beneficial to introduce abstractions for commonly occurring shapes (e.g. an L-shape). By doing this they move towards a more compact language, thereby leading to more efficient communication. Creating such abstractions is a key feature of human collaboration. In this work, we develop a computational framework of artificial agents to solve this kind of collaborative task (Figure 1). Previous work provides computational models of this process [McCarthy et al., 2021, Jergéus et al., 2022], but rely on simplifying assumptions, limiting the ability to capture the collaborative dynamics of language learning, which we address in this work.
We propose a novel multi-agent neuro-symbolic method called Procedural Abstractions for Communicating Efficiently (PACE). We integrate both neural and symbolic learning methods from the computer science literature. On the symbolic side, library learning [Ellis et al., 2021, Bowers et al., 2023], a method from program synthesis that aims to abstract common subprograms into new, more easily reusable terms. On the neural side, emergent communication (EC) [Foerster et al., 2016, Lazaridou and Baroni, 2020], and reinforcement learning (RL) [Sutton and Barto, 1998], enable the development of a flexible, learnable communication language. PACE offers a unified framework for studying the formation and evolution of abstractions across multi-round interactions.
In previous work, EC has been used to explain how communicative pressures for efficient communication shape the language structure in other settings, such as Carlsson et al. [2021, 2024]. Here, we apply EC techniques to study how pressures manifest within abstraction learning.
@@
We evaluate PACE on the Architect-Builder game, which we extend to artificial agents. In our extension, the architect composes programs to describe the goal-scene in an artificial symbolic language (Figure 1 a). The programs are neurally encoded as messages, following conventions from EC, and are then interpreted by the builder, who tries to reconstruct the scene. After several rounds, the architect’s internal symbolic language is extended with a new abstraction for a commonly occurring subprogram (Figure 1 b). Each new abstraction provides alternative shorter ways to express goal-scenes, but require the agents to learn how to communicate and understand them. This contention between shorter but unestablished programs and more verbose established programs is handled via RL over repeated rounds of interaction (Figure 1 c). We find that it exhibits similar tendencies to humans — a development toward a richer language that allows for more concise utterances. Interestingly, after a number of abstractions have been introduced, our model naturally converges to a stable language, after which no more abstractions are introduced. Moreover, we show that in this setting languages that are closer to optimality in terms of the trade-off between average morphosyntactic complexity and language size are easier to learn, connecting our work with the Efficient Communication literature. Our approach addresses limitations of existing approaches and provides a valuable framework for future exploration in this area.
@@
We extend the dataset from McCarthy et al. [2021] for our experiments. We increase the size going from $3$ to $31$ unique shapes, with multiple sub-shapes reoccurring in different shapes. Our shapes are of different sizes and resemble either uppercase or lowercase letters from the English alphabet. As before, the dataset consists of scenes composed of two shapes placed side by side. Our dataset contains $961$ goal-scenes (compared to the original $9$), which enables us to use it for training neural agents. We split the dataset into 930 training scenes and 31 test scenes. These splits are constructed to ensure the distribution over shapes is the same. However, in the test set, the ordered pair of shapes constituting each goal-scene, does not appear in the training set.
@@
There are three phases to PACE, Selection, Communication and Abstraction (see Figure 2). The architect must first choose a program to communicate. For this purpose the architect is initialised with a table of programs written in ${\cal A}_{init}$ for each scene in the training set. Initially there is one program per scene. Secondly, the architect learns how to communicate (neurally) instructions to the builder via EC, allowing them to learn a language for collaboration. After some rounds, the architect enters the Abstraction phase, where it uses a library learning mechanism to identify common sub-sequences to abstract. The new abstraction is then used to add additional (shorter) programs for the relevant scenes. In the next Communication phase, the architect thus has multiple programs to choose from. We use reinforcement learning techniques to learn to choose between alternative programs. We next describe these three phases in more detail. We also provide more detailed algorithms of PACE in the Appendix.
@@
As in other settings like recursive numeral systems [Denić and Szymanik, 2024], programs in the Architect-Builder game can be perfectly informative even with a minimal vocabulary – precisely of size two, with one term referring to the horizontal primitive block and the other referring to the vertical one. With just the two initial action-words the architect can describe infinitely complex goal-scenes, given enough time! This is because the semantics of the environment are compositional. Thus, as Denić and Szymanik [2024] show for recursive numeral systems, we expect that the pressures in competition to shape the agents’ languages in the Architect-Builder game are the average morphosyntactic complexity of the programs describing the goal-scenes and the language size. In our case this corresponds to a trade-off between the average length of the programs communicated by the architect, and the number of lexicalised terms, i.e. the number of primitives plus abstractions in $\mathcal{A}$, the architect’s lexicon.
Implementation Details
The neural agents are deep neural networks which have $1$ and $2$ hidden layers of size $200$, respectively. The architect’s output layer is of size $30$, its maximum vocabulary size. While the output layer of the builder is of size $81$, the size of the grid. We use the gumbel-softmax relaxation with the hard parameter so that the networks can be differentiated end-to-end. We use a learning rate of $0.0009$ in conjunction with the ADAM optimiser [Kingma and Ba, 2014]. The bandit hyperparameters $\alpha$, $\gamma$, $\epsilon$ and $q_{init}$ are set to $0.5$, $0.99$, $0.1$, and $0.0$ respectively. Note, that before an abstraction is introduced we prune all but the $3$ best programs (determined empirically) for each goal-scene in the architects symbolic table to not incur an exponential growth in the number of programs for each goal-scene. We also set the number of epochs $e$ to 40. All hyperparameters chosen are experimentally determined through grid search.
@@
Through empirical experimentation we want to investigate the following questions: 1) Does PACE display conversational tendencies which are similar to humans, allowing for more concise communication after multiple interactions? 2) Does communicative pressures impact what abstractions get adopted? If so, which abstractions are adopted and which are discarded? The behaviour of PACE is compared to two naïve baselines, to assess the impact of each of PACE components better. The first, No abstractions, is PACE without the abstraction phase. The second, Greedy, always picks the shortest program, eliminating the bandit. We use reward (also referred to as reconstruction accuracy) and average morphosyntactic complexity to compare these approaches. Unless otherwise stated results obtained are averaged over $16$ runs.
@@
Figure 7: Two goal-scenes with representations of their initial and final programs chosen by PACE.
@@
McCarthy et al. [2021] propose a Bayesian model of procedural abstractions that combines library learning [Ellis et al., 2021] with social reasoning-based communication [Goodman and Frank, 2016]. While effective in capturing human-like trends, this model assumes a predefined mapping from instructions to builder’s actions, essentially defining a priori the meaning of messages that might not have been introduced yet. They also explicitly constrain language size to prevent adding further abstractions after a certain number have been introduced. Jergéus et al. [2022] take a Deep Reinforcement Learning approach but oversimplify interactions by assuming the builder immediately understands abstractions, collapsing the setup into a single-agent framework and removing pressures for efficient communication. By using EC within PACE, our approach enables both agents to learn interactively, introducing abstractions organically and letting communicative pressures naturally limit language size.
@@ Conclusion
In collaborative task-orientated communication humans tend towards more concise utterances by introducing procedural abstractions. In this work we propose a novel neuro-symbolic algorithm called PACE which displays similar tendencies. Our work serves as a bridge between procedural abstraction learning and efficient communication [Kemp and Regier, 2012, Gibson et al., 2017, Zaslavsky et al., 2019, Gibson et al., 2019, Denić and Szymanik, 2024, Carlsson et al., 2024]. We demonstrate that more optimal languages are easier to learn reinforcing ideas relating learnability with efficient communication.
In future work, we intend to extend our analysis to consider other collaborative domains and deepen the comparison to human behaviour. We also intend to explore how Large Language Models understand and reason about these conversational dynamics. Significant research effort is being invested into exploring the role of natural language as a mechanism for humans to provide instructions to intelligent agents [Brohan et al., 2023, Shi et al., 2024]. Providing these systems with the capability of handling novel procedural abstractions can facilitate improved cooperation between humans and agents. We believe that PACE represents a step towards equipping intelligent agents with flexible and extendable languages.
@@ Acknowledgments
This work was supported by funding from from the Wal- lenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foun- dation and the Swedish Research Council. The computa- tions in this work were enabled by resources provided by the Swedish National Infrastructure for Computing (SNIC). We also extend our thanks to Emil Carlsson and Vikas Garg for insightful discussions at various points throughout the project and to Sandro Stucki for valuable comments on a previous version of the manuscript that lead to this final one.
@@
Dataset Composition
Each goal-scene in the dataset is composed of two shapes placed next to each other, as in Figure 1. These are sampled from 31 shapes we design and report in Figure 10. Each shape is meant to resemble an uppercase or lowercase letter from the English alphabet. We note that there are multiple subshapes that recur several times within the 31 shapes, making our dataset a prime candidate for studying the introduction of abstraction.